# AWM[Uncaptioned image]: Accurate Weight-Matrix Fingerprint for Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) have become foundational to many artificial intelligence applications. However, training an LLM from scratch demands substantial computational resources and vast amounts of data. Consequently, most open-source models are released under specific licenses (Touvron et al., 2023a; Team et al., 2025; Mesnard et al., 2024; Kamath et al., 2025) or require an application (Touvron et al., 2023b; Zhang et al., 2022; Penedo et al., 2023; BaiChuan-Inc, 2023; Team, 2023; Zheng et al., 2023; Grattafiori et al., 2024) and approval process to protect intellectual property. Despite these measures, some developers may circumvent such protections by wrapping or post-training existing base LLMs, then falsely claiming to have trained their own models. Recent controversies (pzc163 et al., 2024; Yoon et al., 2025) have underscored the urgent need to determine whether a suspect model is genuinely trained from scratch or derived from an existing base model.
The core challenge lies in extracting a stable fingerprint to identify the true base model. This task is complicated by the fact that LLMs often undergo heavy post-training processes, such as supervised fine-tuning (SFT), extensive continued pretraining (Team et al., 2024; Hui et al., 2024), and reinforcement learning (RL). Furthermore, emerging techniques like extension to multimodal tasks, architectural pruning (Xia et al., ), and upcycling (He et al., 2024) can drastically alter a model’s parameters, outputs, and even its structure, posing substantial challenges for identification. Moreover, malicious actors might intentionally manipulate weight matrices through operations including scaling, permutation, pruning, and even rotation to obscure a model’s origin.
@@
A review of previous work reveals that existing methods often fail to meet one or more of these criteria. Watermarking techniques, for instance, embed identifiable signals via extra training (Peng et al., 2023; Xu et al., 2024), but this can degrade performance (Russinovich & Salem, 2024), may not survive aggressive post-training (Fernandez et al., 2024; Gubri et al., 2024), and must be applied before the model’s release. Other existing fingerprinting methods either suffer from high false-positive rates (Zhang et al., 2025) or lack robustness against heavy post-training modifications like continued pretraining (Zeng et al., 2024), a limitation we confirm in our experiments.
@@ 2 Related Work
Watermarking. Watermarking methods typically involve finetuning models to inject backdoor triggers that prompt the model to generate predefined content, or embedding watermarks directly into model weights for identification purposes. A substantial body of research has explored watermarking for smaller DNNs such as CNNs and BERT (Devlin et al., 2019), including encoding watermarks into model weights (Chen et al., 2019a; Wang & Kerschbaum, 2021; Liu et al., 2021; Uchida et al., 2017) and injecting triggers to produce specific outputs (Adi et al., 2018; Guo & Potkonjak, 2018; Le Merrer et al., 2020; Chen et al., 2019b). However, these methods are often task-specific and not well-suited for foundation LLMs. For watermarking LLMs, researchers have proposed various methods to inject watermarks for identification (Russinovich & Salem, 2024; Xu et al., 2024; Li et al., 2023; Peng et al., 2023; Kirchenbauer et al., 2023; Zhao et al., 2023). Nevertheless, these approaches inevitably compromise LLM performance and are not robust to extensive post-training.
Fingerprinting. Fingerprinting methods extract intrinsic model features as fingerprints for identification, without requiring additional training and thus without impacting model performance. For small DNNs, prior works (Zhao et al., 2020; Pan et al., 2022; Yang et al., 2022; Lukas et al., 2019; Peng et al., 2022) analyze model behaviors on preset test cases to construct fingerprints. For LLMs, HuRef (Zeng et al., 2024) is a classic weight-based fingerprinting method that derives invariant terms and computes similarities between them for model identification. However, it is not robust to extensive training. REEF (Zhang et al., 2025), from the perspective of representation, measures similarities between model representations, but suffers from a high false positive rate, which limits its practical use. To date, there is still a lack of fingerprinting methods for LLMs that are both robust to extensive training and exhibit a very low risk of false positives.
@@ 3 Preliminary
Most Large Language Models (LLMs) follows a decoder-only Transformer architecture (Radford et al., 2018). The details of various LLMs may differ, yet the Transformer blocks are similar. In particular, a Transformer block in an LLM usually consists of residual connections (He et al., 2016), Root Mean Square Normalization (RMSNorm, Zhang & Sennrich (2019)), the self-attention mechanism (Lin et al., 2017), Rotary Position Embedding (RoPE, Su et al. (2024)), and a feed-forward network (FFN). We denote by $\mathbb{W}_{A}$ the set of an $L$-layered LLM A’s weights, and $\mathbb{W}_{A,\text{partial}}=\{W_{A,\text{emb}}\}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{Q},\text{K}\}\}$ the set of word embeddings and $\text{Q},\text{K}$ matrices, where $W_{A,\text{emb}}$ is the word embeddings and $W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)}$ are the $\text{Q},\text{K}$ matrices at the $l$-th layer. We defer the rest of the notations to Appendix A.1.
@@
Proposed in Kornblith et al. (2019), CKA is a similarity detection method (Zhang et al., 2025) based on Hilbert-Schmidt Independence Criterion (HSIC, Gretton et al. (2005)). It is invariant to column-wise orthogonal transformations and constant multiplications:
@@
Even with semi-orthogonal $U_{1}$ and $U_{2}$ where $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}^{\prime}}$, $U_{2}\in\mathbb{R}^{n_{2}\times n_{2}^{\prime}}$, $n_{1}>n_{1}^{\prime}$, $n_{2}>n_{2}^{\prime}$, CKA still preserves input similarity to some extent (Kang et al., 2025; Chun et al., 2025). Nevertheless, the standard HSIC estimator converges at rate $1/\sqrt{m}$ and shows finite-sample bias (Gretton et al., 2005; Murphy et al., 2024). To this end, an unbiased version (Song et al., 2007) is proposed and extends the value of CKA from $[0,1]$ to [1,1][-1,1]. Additionally, linear kernels are often selected in CKA due to their computational efficiency and similar performance to other kernels (Kornblith et al., 2019). We give the formal definition of CKA in Appendix A.2, and the proof of Theorem 3.1 in Appendix B.1.
@@ 4.1 Problem Definition
Let A and B be two open-source LLMs with the same number of layers $L$, and $\mathbb{W}_{A},\mathbb{W}_{B}$ denote the sets of matrix weights for A and B respectively. Then, if B manipulates A, the manipulations on $W_{A,\text{partial}}$ are
@@
We omit the row-wise transform on word embeddings, $L_{B,\text{emb}}$, since each row of $W_{A,\text{emb}}$ represents a token and mixed token representations are hard to be faithfully recovered in the calculation of attention scores. We further assume that the suspicious and manipulated models produce similar (or identical) outputs to be consistent with the goal of reusing base model performance. This assumption induces constraints on the transformation matrices and, in turn, enables detection via weight-matrix similarity tests. We first develop a fine-grained view of the constraints on $W_{A,\text{partial}}$ in what follows.
@@ 4.2 Residual Connections: Passing Manipulations Forward
A Transformer block consists of multiple components linked by residual connections. If one component takes input $X$, the next receives $Y=X+f(X)$, so any manipulation $T$ on $X$ must be recovered within the component to propagate as the next input (Zeng et al., 2024). This propagation also depends on the constituent functions of the component, which either commute with the manipulation to recover it or remain invariant and absorb it. We formalize this in the following:
Proposition 4.2 (Proof in Appendix B.2).
@@
Compared to Zeng et al. (2024), Section 4.2 shows that residual connections are more vulnerable under a component-wise view, since constituent functions can propagate manipulations in various ways. However, nonlinearities of the functions, especially those within the self-attention mechanism, pose constraints for manipulations on both inputs and weights. We next split the self-attention mechanism into two constituent functions, RMSNorm and attention (see Section A.1), and show how these effects arise and are constrained these two functions with a focus on RMSNorm, RoPE and attention scores.
@@ 4.3 RMSNorm: A Constraint on Embedding Manipulations
Any input to the self-attention mechanism, including word embeddings and hidden states, first go through RMSNorm. However, RMSNorm can facilitate potential manipulations on inputs, because it commutes with certain transformations $R_{B,\text{emb}}$ of the embeddings:
@@
Let models $A$ and $B$ share the same [^1] The conclusions generalize to models with different number of layers (possibly a result of pruning). See Appendix B.3architecture. Let $c\neq 0$ be a scalar, $P$ be a (partial) permutation matrix, and $D$ be a signature matrix. Then, $R_{B,\text{emb}}=cPD$ can be recovered after RMSNorm in model B if related RMSNorm parameters are adjusted.
Theorem 4.3 indicates that RMSNorm is susceptible to embedding manipulations composed of constant multiplications, permutations and sign flips. On the other hand, other manipulations on word embeddings are generally neither commutative with RMSNorm nor invariant to it, which potentially brings a constraint to manipulations. We provide a proof for Theorem 4.3 in Appendix B.3, along with a discussion on other manipulations on word embeddings.
4.4 RoPE and Attention Scores: Boundaries for Q/K Manipulations
After RMSNorm, inputs are fed into attention score calculations. The nonlinearity here, particularly in the softmax and RoPE functions, poses a barrier for manipulations on inputs to pass through. Hence, we follow Zeng et al. (2024) to assume that input manipulations does not change the value of attention scores. The manipulations on Q,K matrices are thus constrained under this assumption.
Theorem 4.4.
The manipulations on Q,K matrices at layer $l$ can be categorized into
  1. 1.
  2. 2.
A proof for Theorem 4.4 is provided in Appendix B.4, where we also show how manipulations on inputs are recovered after V, O matrices to satisfy Section 4.2. These manipulations preserve the attention score values of model A in the suspicious model B. However, they can greatly change the weights of the original model A, bringing difficulty to the development of detection methods.
4.5 Potential Attacks
Combining Theorem 4.3 and Theorem 4.4, the admissible manipulations on $W_{A,\text{partial}}$ (word embeddings and $\text{Q},\text{K}$ matrices) in Section 4.1 are restricted to
$W_{B,\text{emb}}=c\,W_{A,\text{emb}}\,PD+E_{B,\text{emb}}.$ (5)
Here $E$ collects post-training, continual pre-training, pruning, upcycling, multimodal adaptation, or related adjustments. The nonlinear usage of $W_{A,\text{partial}}$, especially through the $\text{Q},\text{K}$ matrices, substantially limits manipulation complexity. Consequently, checking similarity between $W_{A,\text{partial}}$ and $W_{B,\text{partial}}$ is typically sufficient for detection. The remaining avenues targeting $\mathbb{W_{A}}/\mathbb{W}_{A,\text{partial}}$ are deferred to Appendix B.5, where we also discuss how such transformations can recover Transformer block outputs in light of Section 4.2.
5 Methodology
Building on Theorem 4.3 and Theorem 4.4, we propose a two-stage procedure in 1: (i) extract $P$ and $D$ from the word embeddings; (ii) assess cross-model similarity by comparing the Q and K matrices. This design achieves fast, reliable discrimination with minimal computation.
@@ 5 Methodology
A key property of $R_{B,\text{emb}}$ is that the permutation and signature matrices may appear in either order: both $R_{B,\text{emb}}=cPD$ and $R_{B,\text{emb}}=cDP$ are possible. However, any product $DP$ can be rewritten as $P^{\prime}D^{\prime}$ for some permutations $P^{\prime}$ and signature matrices $D^{\prime}$. Hence it suffices to recover a canonical $PD$ from the embeddings. We cast this as a Linear Assignment Problem (LAP; Burkard & Cela (1999)) solved by the Hungarian algorithm (Kuhn, 1955), which is invariant to any nonzero scalar factor. Specifically, we first restrict to the shared vocabulary and form the matrix of absolute cosine similarities between the columns of $W_{A,\text{emb}}$ and $W_{B,\text{emb}}$. Then, LAP is applied to obtain the permutation $P$. Next, we use the signs of the cosine similarities at the matched pairs to reconstruct the signature matrix $D$. This approach effectively reconstruct corresponding transformations due to its low demand of additional parameters in the detection process.
@@
Despite accounting for permutation and signature manipulations, the orthogonal transformations $U_{B,i}^{(l)}$ remain challenging. They introduce a substantial nuisance parameter burden: if $W_{A,i}^{(l)}\in\mathbb{R}^{d\times n}$, then orthogonal $U_{B,i}^{(l)}\in\mathbb{R}^{d\times d}$ contributes $d^{2}$ parameters, which is prohibitive when $d\approx n$ and can undermine robustness by adding parameters to detections (Simmons et al., 2011). We therefore use Central Kernel Alignment (CKA) as a parameter-free similarity metric: by Theorem 3.1, CKA is invariant to orthogonal transforms and constant rescaling. While this invariance does not extend to semi-orthogonal $U_{B,i}^{(l)}$ (e.g., induced by pruning), CKA remains effective to a meaningful extent in such cases, relieving the need to explicitly reconstruct $U_{B,i}^{(l)}$. To further mitigate biases (Gretton et al., 2005; Murphy et al., 2024), we adopt the unbiased variant (Song et al., 2007), termed UCKA (see Appendix A.2).

### v2

@@ 1 Introduction
Large language models (LLMs) have become foundational to many artificial intelligence applications. However, training an LLM from scratch demands substantial computational resources and vast amounts of data. Consequently, most open-source models are released under specific licenses (touvron2023llama; team2025kimi; team2024gemma; team2025gemma) or require an application (touvron2023llama2; zhang2022opt; falcon; baichuan; 2023internlm; zheng2023codegeex; grattafiori2024llama) and approval process to protect intellectual property. Despite these measures, some developers may circumvent such protections by wrapping or post-training existing base LLMs, then falsely claiming to have trained their own models. Recent controversies (pzc1632024llama3v; yoon2025intrinsic) have underscored the urgent need to determine whether a suspect model is genuinely trained from scratch or derived from an existing base model.
The core challenge lies in extracting a stable fingerprint to identify the true base model. This task is complicated by the fact that LLMs often undergo heavy post-training processes, such as supervised fine-tuning (SFT), extensive continued pretraining (team2024codegemma; hui2024qwen2), and reinforcement learning (RL). Furthermore, emerging techniques like extension to multimodal tasks, architectural pruning (xiasheared), and upcycling (he2024upcycling) can drastically alter a model’s parameters, outputs, and even its structure, posing substantial challenges for identification. Moreover, malicious actors might intentionally manipulate weight matrices through operations including scaling, permutation, pruning, and even rotation to obscure a model’s origin.
@@
A review of previous work reveals that existing methods often fail to meet one or more of these criteria. Watermarking techniques, for instance, embed identifiable signals via extra training (peng2023you; xu2024instructional), but this can degrade performance (russinovich2024hey), may not survive aggressive post-training (fernandez2024functional; gubri2024trap), and must be applied before the model’s release. Other existing fingerprinting methods either suffer from high false-positive rates (zhang2025reef) or lack robustness against heavy post-training modifications like continued pretraining (zeng2024huref), a limitation we confirm in our experiments.
@@ 2 Related Work
Watermarking. Watermarking methods typically involve finetuning models to inject backdoor triggers that prompt the model to generate predefined content, or embedding watermarks directly into model weights for identification purposes. A substantial body of research has explored watermarking for smaller DNNs such as CNNs and BERT (devlin2019bert), including encoding watermarks into model weights (chen2019deepmarks; white_wang2021riga; white_liu2021residuals; Uchida_2017) and injecting triggers to produce specific outputs (black_adi2018turning; black_guo2018watermarking; black_le2020adversarial; black_chen2019blackmarks). However, these methods are often task-specific and not well-suited for foundation LLMs. For watermarking LLMs, researchers have proposed various methods to inject watermarks for identification (russinovich2024hey; xu2024instructional; li2023plmmark; peng2023you; kirchenbauer2023watermark; zhao2023protecting). Nevertheless, these approaches inevitably compromise LLM performance and are not robust to extensive post-training.
Fingerprinting. Fingerprinting methods extract intrinsic model features as signatures for identification without requiring additional training, thereby preserving model performance. These methods are generally categorized based on the auditor’s access level: white-box (full access) and black-box (API access).
White-box Fingerprinting. When model parameters are accessible, auditors can derive fingerprints from static weights or dynamic internal states. For small DNNs, prior works (fingerprinting_zhao2020afa; fingerprinting_pan2022metav; fingerprinting_yang2022metafinger; lukas2019deep; fingerprinting_peng2022uapfp) typically analyze model behaviors on preset test cases. In the context of LLMs, HuRef (zeng2024huref) is a representative static method that derives invariant terms from weight matrices to compute similarity. However, it is not robust to extensive training. Similarly, yoon2025intrinsic utilize the standard deviation distributions of attention matrices as fingerprints; while stable under training, such statistical measures may carry a risk of false positives. Moving beyond static weights, other methods investigate dynamic signals. For instance, DeepJudge (chen2022copy) and EasyDetector (zhang2024easydetector) utilize intermediate activation values to quantify model distance. REEF (zhang2025reef) further measures the geometric similarity of representation spaces but also suffers from a high false positive rate. More recently, TensorGuard (wu2025gradient) proposes utilizing the statistical features of gradients generated during backpropagation as a stable signature to characterize the model’s optimization landscape.
Black-box Fingerprinting. With the proliferation of Model-as-a-Service (MaaS), black-box methods that rely solely on API input-output interactions have gained traction. These approaches generally follow two paradigms: analyzing output distributions or constructing specific trigger queries. The former focuses on identifying unique stylistic idiosyncrasies or probability distributions inherent to a model family. For example, LLMmap (pasquini2025llmmap) and other distributional approaches (yang2024fingerprint; gao2025model; mcgovern2025your) query the model with general prompts to capture distinct response patterns or logit features. The latter paradigm involves optimizing specific ”trap” inputs or adversarial queries (xu2024instructional; gubri2024trap; jin2024proflingo; xu2025rap) designed to force a suspect model to output a predefined unique response.
However, black-box methods face significant robustness challenges. Fundamentally, relying solely on surface-level outputs results in a loss of critical information regarding the model’s internal mechanisms compared to white-box access (shao2025reading). Consequently, these methods are often fragile to post-training modifications such as instruction tuning or simple system prompt changes, which can disrupt the fingerprint (xu2025rap; tsai2025rofl). To date, there is still a lack of fingerprinting methods for LLMs that are both robust to extensive training and exhibit a very low risk of false positives.
@@ 3 Preliminary
Most Large Language Models (LLMs) follows a decoder-only Transformer architecture (radford2018improving). The details of various LLMs may differ, yet the Transformer blocks are similar. In particular, a Transformer block in an LLM usually consists of residual connections (he2016deep), Root Mean Square Normalization (RMSNorm, zhang2019root), the self-attention mechanism (lin2017structured), Rotary Position Embedding (RoPE, su2024roformer), and a feed-forward network (FFN). We denote by $\mathbb{W}_{A}$ the set of an $L$-layered LLM A’s weights, and $\mathbb{W}_{A,\text{partial}}=\{W_{A,\text{emb}}\}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{Q},\text{K}\}\}$ the set of word embeddings and $\text{Q},\text{K}$ matrices, where $W_{A,\text{emb}}$ is the word embeddings and $W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)}$ are the $\text{Q},\text{K}$ matrices at the $l$-th layer. We defer the rest of the notations to Appendix D.1.
@@
Proposed in kornblith2019similarity, CKA is a similarity detection method (zhang2025reef) based on Hilbert-Schmidt Independence Criterion (HSIC, gretton2005measuring). It is invariant to column-wise orthogonal transformations and constant multiplications:
@@
Even with semi-orthogonal $U_{1}$ and $U_{2}$ where $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}^{\prime}}$, $U_{2}\in\mathbb{R}^{n_{2}\times n_{2}^{\prime}}$, $n_{1}>n_{1}^{\prime}$, $n_{2}>n_{2}^{\prime}$, CKA still preserves input similarity to some extent (kang2025spectral; chun2025estimating). Nevertheless, the standard HSIC estimator converges at rate $1/\sqrt{m}$ and shows finite-sample bias (gretton2005measuring; murphy2024correcting). To this end, an unbiased version (song2007supervised) is proposed and extends the value of CKA from $[0,1]$ to [1,1][-1,1]. Additionally, linear kernels are often selected in CKA due to their computational efficiency and similar performance to other kernels (kornblith2019similarity). We give the formal definition of CKA in Appendix D.2, and the proof of Theorem 3.1 in Appendix E.1.
@@ 4 Manipulations on An Open-Source LLM
In this part, we investigate which matrix-weight manipulations can remain compatible with preserving a model’s behaviour. We first present the definition of matrix weight manipulation in Section˜4.1. Then, we examine how key components of an LLM, including residual connections (Section˜4.2), RMSNorm (Section˜4.3), and RoPE together with attention scores (Section˜4.4), can constrain and gradually narrow the space of admissible manipulations when the manipulated model is required to produce outputs similar to the base model. Finally, we derive potential attacks on matrix weights in Section˜4.5 based on these constraints, with a focus on the Q, K matrices and the embedding layer.
@@ 4.1 Problem Definition
Let A and B be two open-source LLMs with the same number of layers $L$, and $\mathbb{W}_{A},\mathbb{W}_{B}$ denote the sets of matrix weights for A and B respectively. Then, if B manipulates A, the manipulations on $\mathbb{W}_{A,\text{partial}}$ are
@@
We omit the row-wise transform on word embeddings, $L_{B,\text{emb}}$, since each row of $W_{A,\text{emb}}$ represents a token and mixed token representations are hard to be faithfully recovered in the calculation of attention scores. We further assume that the suspicious and manipulated models produce similar (or identical) outputs to be consistent with the goal of reusing base model performance. This assumption induces constraints on the transformation matrices and, in turn, enables detection via weight-matrix similarity tests. We first develop a fine-grained view of the constraints on $\mathbb{W}_{A,\text{partial}}$ in what follows.
@@ 4.2 Residual Connections: Passing Manipulations Forward
In what follows, we first analyze how residual connections propagate weight manipulations forward through the network while preserving the model’s outputs. A Transformer block consists of multiple components linked by residual connections. If one component takes input $X$, the next receives $Y=X+f(X)$, so any manipulation $T$ on $X$ must be recovered within the component to propagate as the next input (zeng2024huref). This propagation also depends on the constituent functions of the component, which either commute with the manipulation to recover it or remain invariant and absorb it. We formalize this in the following:
Proposition 4.2 (Proof in Appendix E.2).
@@
Compared to zeng2024huref, Section 4.2 shows that residual connections are more vulnerable under a component-wise view, since constituent functions can propagate manipulations in various ways. However, nonlinearities of the functions, especially those within the self-attention mechanism, pose constraints for manipulations on both inputs and weights. We next split the self-attention mechanism into two constituent functions, RMSNorm and attention (see Section D.1), and show how these effects arise and are constrained these two functions with a focus on RMSNorm, RoPE and attention scores.
@@ 4.3 RMSNorm: A Constraint on Embedding Manipulations
Next, we examine how RMSNorm may allow certain manipulations of word embeddings. Any input to the self-attention mechanism, including word embeddings and hidden states, first go through RMSNorm. However, RMSNorm can facilitate potential manipulations on inputs, because it commutes with certain transformations $R_{B,\text{emb}}$ of the embeddings:
@@
Let models $A$ and $B$ share the same [^1] The conclusions generalize to models with different number of layers (possibly a result of pruning). See Appendix E.3architecture. Let $c\neq 0$ be a scalar, $P$ be a (partial) permutation matrix, and $D$ be a signature matrix. Then, $R_{B,\text{emb}}=cPD$ can be recovered after RMSNorm in model B if related RMSNorm parameters are adjusted.
Theorem 4.3 indicates that RMSNorm is susceptible to embedding manipulations composed of constant multiplications, permutations and sign flips. On the other hand, other manipulations on word embeddings are generally neither commutative with RMSNorm nor invariant to it, which potentially brings a constraint to manipulations. We provide a proof for Theorem 4.3 in Appendix E.3, along with a discussion on other manipulations on word embeddings.
@@
4.4 RoPE and Attention Scores: Boundaries for Q/K Manipulations
Finally, we study RoPE and the attention score computation to characterize which manipulations on Q,K matrices can preserve attention scores. After RMSNorm, inputs are fed into attention score calculations. The nonlinearity here, particularly in the softmax and RoPE functions, poses a barrier for manipulations on inputs to pass through. Hence, we follow zeng2024huref to assume that input manipulations does not change the value of attention scores. The manipulations on Q,K matrices are thus constrained under this assumption.
Theorem 4.4.
The manipulations on Q,K matrices at layer $l$ can be categorized into
  1. 1.
  2. 2.
A proof for Theorem 4.4 is provided in Appendix E.4, where we also show how manipulations on inputs are recovered after V, O matrices to satisfy Section 4.2. These manipulations preserve the attention score values of model A in the suspicious model B. However, they can greatly change the weights of the original model A, bringing difficulty to the development of detection methods.
4.5 Potential Attacks
Combining Theorem 4.3 and Theorem 4.4, the admissible manipulations on $\mathbb{W}_{A,\text{partial}}$ (word embeddings and $\text{Q},\text{K}$ matrices) in Section 4.1 are restricted to
$W_{B,\text{emb}}=c\,W_{A,\text{emb}}\,PD+E_{B,\text{emb}}.$ (5)
Here $E$ collects post-training, continual pre-training, pruning, upcycling, multimodal adaptation, or related adjustments. The nonlinear usage of $\mathbb{W}_{A,\text{partial}}$, especially through the $\text{Q},\text{K}$ matrices, substantially limits manipulation complexity. Consequently, checking similarity between $\mathbb{W}_{A,\text{partial}}$ and $\mathbb{W}_{B,\text{partial}}$ is typically sufficient for detection. The remaining avenues targeting $\mathbb{W_{A}}/\mathbb{W}_{A,\text{partial}}$ are deferred to Appendix E.5, where we also discuss how such transformations can recover Transformer block outputs in light of Section 4.2.
5 Methodology
Building on Theorem 4.3 and Theorem 4.4, we propose a two-stage procedure in Algorithm 1: (i) extract $P$ and $D$ from the word embeddings; (ii) assess cross-model similarity by comparing the Q and K matrices. This design achieves fast, reliable discrimination with minimal computation.
@@ 5 Methodology
A key property of $R_{B,\text{emb}}$ is that the permutation and signature matrices may appear in either order: both $R_{B,\text{emb}}=cPD$ and $R_{B,\text{emb}}=cDP$ are possible. However, any product $DP$ can be rewritten as $P^{\prime}D^{\prime}$ for some permutations $P^{\prime}$ and signature matrices $D^{\prime}$. Hence it suffices to recover a canonical $PD$ from the embeddings. We cast this as a Linear Assignment Problem (LAP; burkard1999linear) solved by the Hungarian algorithm (kuhn1955hungarian), which is invariant to any nonzero scalar factor. Specifically, we first restrict to the shared vocabulary and form the matrix of absolute cosine similarities between the columns of $W_{A,\text{emb}}$ and $W_{B,\text{emb}}$. Then, LAP is applied to obtain the permutation $P$. Next, we use the signs of the cosine similarities at the matched pairs to reconstruct the signature matrix $D$. This approach effectively reconstruct corresponding transformations due to its low demand of additional parameters in the detection process.
@@
Despite accounting for permutation and signature manipulations, the orthogonal transformations $U_{B,i}^{(l)}$ remain challenging. They introduce a substantial nuisance parameter burden: if $W_{A,i}^{(l)}\in\mathbb{R}^{d\times n}$, then orthogonal $U_{B,i}^{(l)}\in\mathbb{R}^{d\times d}$ contributes $d^{2}$ parameters, which is prohibitive when $d\approx n$ and can undermine robustness by adding parameters to detections (simmons2011false). We therefore use Central Kernel Alignment (CKA) as a parameter-free similarity metric: by Theorem 3.1, CKA is invariant to orthogonal transforms and constant rescaling. While this invariance does not extend to semi-orthogonal $U_{B,i}^{(l)}$ (e.g., induced by pruning), CKA remains effective to a meaningful extent in such cases, relieving the need to explicitly reconstruct $U_{B,i}^{(l)}$. To further mitigate biases (gretton2005measuring; murphy2024correcting), we adopt the unbiased variant (song2007supervised), termed UCKA (see Appendix D.2).